Aconsumer consuming two commodities x and y has the following utility functions U= xy + 2z. If the price of the two commodities are 4 and 2 respectively and his/her budget is birr 60.





Find the quantities of good x and y which with miximize utility





Find the MRSxy at optimum

1
Expert's answer
2022-03-22T11:08:18-0400

A.


U=XY+2X


budget line


I=PxX+PyY , where Px=4,Py=2 ,Hence


I=4X+2Y


"\frac{MUx}{MUy}"


At maximum Utility


"\frac{MUx}{MUy}=\frac{Px}{Py}"


MUx="\frac{du}{dx}" XY+2X

=Y+2

MUy="\frac{du}{dy}" XY+2X

=X

Now

"\frac{MUx}{MUy}=\frac{Px}{Py}"


"\frac{Y+2}{X}=\frac{4}{2}"


"\frac{Y+2}{X}=2"


Y+2=2X


Y=2X-2 ,substitute this to budget line:

60=4X+2Y

60=4X+2"\times" (2X-2)

60=4X+4X-4

64=8x

x=8


Y=2X-2

Y=(2"\times8" )-2

Y=16-2

Y=14

At maximum utility X=8,Y=14


B).


MRSxy="\frac{MUx}{MUy }"


="\frac{Y+2}{X}" ,where Y=14,X=8


="\frac{14+2}{8}"


="\frac{16}{8}"


=2


Hence, MRSxy=2